# Optimal and fast online change point estimation in linear regression

### v1

@@ 5 Proofs
$\mathbb{P}\left(\max_{1\leq i\leq n}\left|f_{-}\left(\frac{i}{n}\right)-\hat{f}_{-}\left(\frac{i}{n}\right)\right|\geq\epsilon\right)\leq\mathbb{P}\left(\left\{\left|f_{-}(0)-\hat{f}_{-}(0)\right|\geq\epsilon\right\}\cup\left\{\left|f_{-}(1)-\hat{f}_{-}(1)\right|\geq\epsilon\right\}\right)$
=($\qquad=\mathbb{P}\left(\left\{\left|\alpha_{-}-\beta_{-}\tau_{0}-\hat{\alpha}\right|\geq\epsilon\right\}\cup\left\{\left|\alpha_{-}+\beta_{-}\left(1-\tau_{0}\right)-\hat{\alpha}-\hat{\beta}\right|\geq\epsilon\right\}\right)$
($\qquad\leq\mathbb{P}\left(\left\{\left|\alpha_{-}-\beta_{-}\tau_{0}-\hat{\alpha}\right|\geq\epsilon\right\}\cup\left\{\left|\beta_{-}-\hat{\beta}\right|\geq\frac{\epsilon}{2}\right\}\right)$
$\qquad\leq\mathbb{P}\left(\left|\alpha_{-}-\beta_{-}\tau_{0}-\hat{\alpha}\right|\geq\epsilon\right)+\mathbb{P}\left(\left|\beta_{-}-\hat{\beta}\right|\geq\frac{\epsilon}{2}\right)$
$\qquad=\mathbb{P}\left(\left|Z_{\alpha}\right|\geq\epsilon\sqrt{\frac{c^{2}n^{2}-cn}{4cn+2}}\right)+\mathbb{P}\left(\left|Z_{\beta}\right|\geq\frac{\epsilon}{2}\sqrt{\frac{c^{3}n^{2}-c}{12n}}\right)$
$\qquad\leq\exp\left\{-\frac{\epsilon^{2}\left(c^{2}n^{2}-cn\right)}{8cn+4}\right\}+\exp\left\{-\frac{\epsilon^{2}(c^{3}n^{2}-c)}{96n}\right\},$ (10)
@@
$\mathbb{P}\left(\max_{1\leq i\leq n}\left|f_{-}\left(\frac{i}{n}\right)-\hat{f}_{-}\left(\frac{i}{n}\right)\right|\geq\frac{c}{10}\right)$ $\leq\exp\left\{-\frac{c^{2}\left(c^{2}n^{2}-cn\right)}{800cn+4}\right\}+\exp\left\{-\frac{c^{2}(c^{3}n^{2}-c)}{9600n}\right\}$
@@
To summarize, we have shown that under the event $\mathcal{A}$, it holds that $0\leq\hat{\tau}_{J,n}-\tau_{0}\leq 3N/n$ and $\mathbb{P}\left(\mathcal{A}^{c}\right)\leq n^{-4}$. Thus, since $11/n^{2}\log^{2}n<$ 1 for $n\geq 2$,
$\max_{\theta_{0}\in\Theta^{J}}\mathbb{E}_{\tau_{0}}\left[\left(\frac{n}{\log n}\left(\hat{\tau}_{J,n}-\tau_{0}\right)\right)^{2}\right]$
$\qquad=\max_{\theta_{0}\in\Theta^{J}}\left(\mathbb{E}_{\tau_{0}}\left[\left(\frac{n}{\log n}\left(\hat{\tau}_{J,n}-\tau_{0}\right)\right)^{2}\mathbb{I}(\mathcal{A})\right]+\mathbb{E}_{\tau_{0}}\left[\left(\frac{n}{\log n}\left(\hat{\tau}_{J,n}-\tau_{0}\right)\right)^{2}\mathbb{I}(\mathcal{A}^{c})\right]\right)$
$\qquad\leq\max_{\theta_{0}\in\Theta^{J}}\left(\mathbb{E}_{\tau_{0}}\left[\left(\frac{3N}{\log n}\right)^{2}\mathbb{I}(\mathcal{A})\right]+\mathbb{E}_{\tau_{0}}\left[\left(\frac{n}{\log n}\right)^{2}\mathbb{I}(\mathcal{A}^{c})\right]\right)$
$\qquad\leq\left(\frac{3N}{\log n}\right)^{2}+\left(\frac{n}{\log n}\right)^{2}n^{-4}\leq 9b^{2}+1=9\frac{10^{6}}{4c^{4}}+1.$
@@
In the case of a kink, we assume that $\left|\beta_{+}-\beta_{-}\right|\geq\delta_{0}$ and $\left|\alpha_{+}-\alpha_{-}\right|<\delta_{0}$. Here we set $N\coloneqq bn^{2/3}\log^{1/3}n$, assumed to be an integer, with $b$ positive and independent of $n$ and $M\coloneqq 2N+(m\mod N)$. We set the threshold as $\rho_{K}=4c/5n$ for some fixed $c>0$. We scale our test statistic by $d_{M}\coloneqq M(M+1)(2M+1)/6$. For $m<n\tau_{0}$, it holds
@@
And for $m_{0}\coloneqq\left\lceil n\tau_{0}\right\rceil+3N-1$, we get
@@
As there is a kink and $m_{0}-M-n\tau_{0}\geq 0$. we get for the summand (13a), in the case of $\left|\alpha_{+}-\alpha_{-}\right|=0$,
@@
$\qquad=\left|\frac{1}{d_{M}}\sum_{i=1}^{M}i\left(\left(\beta_{+}-\beta_{-}\right)\left(\frac{m_{0}-M+i}{n}-\tau_{0}\right)\right)\right|$
$\qquad\geq\frac{\left|\beta_{+}-\beta_{-}\right|}{d_{M}}\sum_{i=1}^{M}i\left(\frac{m_{0}-M+i}{n}-\tau_{0}\right)\geq\frac{\left|\beta_{+}-\beta_{-}\right|}{d_{M}n}\sum_{i=1}^{M}i^{2}>\frac{c}{n},$
for $\left|\beta_{+}-\beta_{-}\right|>c$. In the case of $\left|\alpha_{+}-\alpha_{-}\right|>0$, as $m_{0}-M-n\tau_{0}\leq N$, the summand (13a) can be bounded as follows
@@
$\qquad=\left|\frac{1}{d_{M}}\sum_{i=1}^{M}i\left(\left(\beta_{+}-\beta_{-}\right)\left(\frac{m_{0}-M+i}{n}-\tau_{0}\right)+\alpha_{+}-\alpha_{-}\right)\right|$
$\qquad\geq\frac{\left|\alpha_{+}-\alpha_{-}\right|}{d_{M}}\sum_{i=1}^{M}i-\frac{\left|\beta_{+}-\beta_{-}\right|}{d_{M}}\sum_{i=1}^{M}i\left(\frac{m_{0}-M+i}{n}-\tau_{0}\right)$
$\qquad\geq\frac{3\left|\alpha_{+}-\alpha_{-}\right|}{2M+1}-\frac{\left|\beta_{+}-\beta_{-}\right|}{d_{M}n}\sum_{i=1}^{M}i(i+N)=\frac{3\left|\alpha_{+}-\alpha_{-}\right|}{2M+1}-\frac{\left|\beta_{+}-\beta_{-}\right|}{n}\left(1+\frac{3N}{2M+1}\right)$
$\qquad>\frac{c}{n},$
for $n$ large enough. Set $Z_{m}\coloneqq(1/\sqrt{d_{M}})\sum_{i=1}^{M}i\varepsilon_{m-M+i}$ for $m=k+1,...,n$ standard normal distributed. Next we define the event
$\mathcal{A}\coloneqq\left\{\max_{k<m\leq n}|Z_{m}|<\sqrt{8\log n},\max_{1\leq i\leq n}|f_{-}\left(\frac{i}{n}\right)-\hat{f}_{-}\left(\frac{i}{n}\right)|<\frac{c\left(2M+1\right)}{30n}\right\}.$
@@
$\qquad=\sqrt{\frac{3\log n}{b^{3}n^{2}\log n}}=\frac{c}{10n}.$
@@
$\left|\frac{1}{d_{M}}\sum_{i=1}^{M}i\left(f_{-}\left(\frac{m-M+i}{n}\right)-\hat{f}_{-}\left(\frac{m-M+i}{n}\right)\right)\right|<\frac{1}{d_{M}}\sum_{i=1}^{M}i\frac{c\left(2M+1\right)}{30n}$
$\qquad=\frac{6}{M\left(M+1\right)\left(2M+1\right)}\frac{c\left(2M+1\right)}{30n}\frac{M\left(M+1\right)}{2}=\frac{c}{10n}.$
@@
$\mathcal{A}\subset\left\{K_{m}<\frac{c}{5n}\text{ for }m<n\tau_{0}\text{ and }K_{m}\geq\frac{4c}{5n}\text{ for }m=\left\lceil n\tau_{0}\right\rceil+3N-1\right\}\subset\left\{n\tau_{0}\leq n\hat{\tau}_{K,n}\leq n\tau_{0}+3N\right\}$
@@
$\mathbb{P}_{\tau_{0}}\left(\max_{k<m\leq n}|Z_{m}|\geq\sqrt{8\log n}\right)\leq n\exp\left\{-4\log n\right\}=n^{-3}.$ (15)
@@
$\mathbb{P}_{\tau_{0}}\left(\max_{1\leq i\leq n}\left|f_{-}\left(\frac{i}{n}\right)-\hat{f}_{-}\left(\frac{i}{n}\right)\right|\geq\frac{c\left(2N+1\right)}{10n}\right)$
$\qquad\leq\exp\left\{-\frac{c^{2}(2N+1)^{2}\left(c^{2}n^{2}-cn\right)}{100n^{2}(8cn+4)}\right\}+\exp\left\{-\frac{c^{2}(2N+1)^{2}(c^{3}n^{2}-c)}{9600n^{3}}\right\},$
$\qquad\asymp\exp\left\{-n^{1/3}\log^{2/3}n\right\}\ll n^{-3}.$
@@
$\mathbb{P}_{\tau_{0}}\left(\mathcal{A}^{c}\right)\leq\mathbb{P}_{\tau_{0}}\left(\max_{1\leq m\leq M}|Z_{m}|\geq\sqrt{8\log M}\right)+\mathbb{P}_{\tau_{0}}\left(\max_{1\leq i\leq n}\left|f_{-}\left(\frac{i}{n}\right)-\hat{f}_{-}\left(\frac{i}{n}\right)\right|\geq\frac{c\left(2N+1\right)}{10n}\right)\leq n^{-3}.$
@@
$\max_{\theta_{0}\in\Theta^{K}}\mathbb{E}_{\tau_{0}}\left[\left(\frac{n^{1/3}}{\log^{1/3}n}\left(\tau_{K,n}-\tau_{0}\right)\right)^{2}\right]$
$\qquad=\max_{\theta_{0}\in\Theta^{K}}\left(\mathbb{E}_{\tau_{0}}\left[\left(\frac{n^{1/3}}{\log^{1/3}n}\left(\tau_{K,n}-\tau_{0}\right)\right)^{2}\mathbb{I}(\mathcal{A})\right]+\mathbb{E}_{\tau_{0}}\left[\left(\frac{n^{1/3}}{\log^{1/3}n}\left(\tau_{K,n}-\tau_{0}\right)\right)^{2}\mathbb{I}(\mathcal{A}^{c})\right]\right)$
$\qquad\leq\max_{\theta_{0}\in\Theta}\left(\mathbb{E}_{\tau_{0}}\left[\left(\frac{3N}{n^{2/3}\log^{1/3}n}\right)^{2}\mathbb{I}(\mathcal{A})\right]+\mathbb{E}_{\tau_{0}}\left[\left(\frac{n^{1/3}}{\log^{1/3}n}\right)^{2}\mathbb{I}(\mathcal{A}^{c})\right]\right)$
$\qquad\leq\left(\frac{3N}{n^{2/3}\log^{1/3}n}\right)^{2}+\left(\frac{n^{1/3}}{\log^{1/3}n}\right)^{2}n^{-3}\leq 9b^{2}+1=9\cdot\left(\frac{300}{c^{2}}\right)^{2/3}+1,$

### v2

@@ 5 Proofs
$\mathbb{P}\left(\max_{1\leq i\leq n}\left|f_{-}\left(\frac{i}{n}\right)-\hat{f}_{-}\left(\frac{i}{n}\right)\right|\geq\epsilon\right)=\mathbb{P}\left(\left\{\left|f_{-}(0)-\hat{f}_{-}(0)\right|\geq\epsilon\right\}\cup\left\{\left|f_{-}(1)-\hat{f}_{-}(1)\right|\geq\epsilon\right\}\right)$
$=\,$ $\mathbb{P}\left(\left\{\left|\alpha_{-}-\beta_{-}\tau-\hat{\alpha}\right|\geq\epsilon\right\}\cup\left\{\left|\alpha_{-}+\beta_{-}\left(1-\tau\right)-\hat{\alpha}-\hat{\beta}\right|\geq\epsilon\right\}\right)$
$\leq\,$ $\mathbb{P}\left(\left\{\left|\alpha_{-}-\beta_{-}\tau-\hat{\alpha}\right|\geq\frac{\epsilon}{2}\right\}\cup\left\{\left|\beta_{-}-\hat{\beta}\right|\geq\frac{\epsilon}{2}\right\}\right)$
$\leq\,$ $\mathbb{P}\left(\left|\alpha_{-}-\beta_{-}\tau-\hat{\alpha}\right|\geq\frac{\epsilon}{2}\right)+\mathbb{P}\left(\left|\beta_{-}-\hat{\beta}\right|\geq\frac{\epsilon}{2}\right)$
$=\,$ $\mathbb{P}\left(\left|Z_{\alpha}\right|\geq\frac{\epsilon}{2}\sqrt{\frac{k^{2}-k}{4k+2}}\right)+\mathbb{P}\left(\left|Z_{\beta}\right|\geq\frac{\epsilon}{2}\sqrt{\frac{k^{3}-k}{12n^{2}}}\right)$
$\leq\,$ $\exp\left\{-\frac{\epsilon^{2}\left(k^{2}-k\right)}{32k+16}\right\}+\exp\left\{-\frac{\epsilon^{2}(k^{3}-k)}{96n^{2}}\right\},$ (15)
@@
$\mathbb{P}\left(\max_{1\leq i\leq n}\left|f_{-}\left(\frac{i}{n}\right)-\hat{f}_{-}\left(\frac{i}{n}\right)\right|\geq\frac{c}{10}\right)$ $\leq\exp\left\{-\frac{c^{2}\left(c^{2}n^{2}-cn\right)}{3200cn+16}\right\}+\exp\left\{-\frac{c^{2}(c^{3}n^{2}-c)}{9600n}\right\}$
@@
To summarize, we have shown that under the event $\mathcal{A}$, it holds that $0\leq\hat{\tau}_{J,n}-\tau\leq 3N/n$ and $\mathbb{P}\left(\mathcal{A}^{c}\right)\leq n^{-4}$. Thus, since $1/\left(n^{2}\log^{2}n\right)<1$ for $n\geq 2$,
$\max_{\theta_{0}\in\Theta^{J}}\mathbb{E}_{\tau}\left[\left(\frac{n}{\log n}\left(\hat{\tau}_{J,n}-\tau\right)\right)^{2}\right]$
$\qquad=\max_{\theta_{0}\in\Theta^{J}}\left(\mathbb{E}_{\tau}\left[\left(\frac{n}{\log n}\left(\hat{\tau}_{J,n}-\tau\right)\right)^{2}\mathbb{I}(\mathcal{A})\right]+\mathbb{E}_{\tau}\left[\left(\frac{n}{\log n}\left(\hat{\tau}_{J,n}-\tau\right)\right)^{2}\mathbb{I}(\mathcal{A}^{c})\right]\right)$
$\qquad\leq\max_{\theta_{0}\in\Theta^{J}}\left(\mathbb{E}_{\tau}\left[\left(\frac{3N}{\log n}\right)^{2}\mathbb{I}(\mathcal{A})\right]+\mathbb{E}_{\tau}\left[\left(\frac{n}{\log n}\right)^{2}\mathbb{I}(\mathcal{A}^{c})\right]\right)$
$\qquad\leq\left(\frac{3N}{\log n}\right)^{2}+\left(\frac{n}{\log n}\right)^{2}n^{-4}\leq 9b^{2}+1=9\cdot\frac{10^{6}}{4c^{4}}+1.$
@@
We additionally note that the choice $k\asymp n$ is required to establish (16). However, by leveraging (15), this requirement can be relaxed to $k\asymp n^{2/3}\log(n)^{1/3}$.
@@
We assume that $\left|\beta_{+}-\beta_{-}\right|\geq\delta_{0}$ and $\left|\alpha_{+}-\alpha_{-}\right|<\delta_{0}$. Note that $N\coloneqq bn^{2/3}\log^{1/3}n$, with $b$ positive and independent of $n$, and $M\coloneqq 2N+(m\mod N)$. The threshold is $\rho_{K}=4c/(5n)$ for some fixed $c>0$. Let $d_{M}\coloneqq M(M+1)(2M+1)/6$. For $m<n\tau$, it holds
@@
And for $m_{0}\coloneqq\left\lceil n\tau\right\rceil+3N-1$, we get
@@
As there is a kink and $m_{0}-M-n\tau\geq 0$. we get for the summand (18a), in the case of $\left|\alpha_{+}-\alpha_{-}\right|=0$,
@@
$\qquad=\left|\frac{1}{d_{M}}\sum_{i=1}^{M}i\left(\left(\beta_{+}-\beta_{-}\right)\left(\frac{m_{0}-M+i}{n}-\tau\right)\right)\right|$
$\qquad\geq\frac{\left|\beta_{+}-\beta_{-}\right|}{d_{M}}\sum_{i=1}^{M}i\left(\frac{m_{0}-M+i}{n}-\tau\right)\geq\frac{\left|\beta_{+}-\beta_{-}\right|}{d_{M}n}\sum_{i=1}^{M}i^{2}>\frac{c}{n},$
for $\left|\beta_{+}-\beta_{-}\right|>c$. In the case of $\left|\alpha_{+}-\alpha_{-}\right|>0$, as $m_{0}-M-n\tau\leq N$, the summand (18a) can be bounded as follows
@@
$=\,$ $\left|\frac{1}{d_{M}}\sum_{i=1}^{M}i\left(\left(\beta_{+}-\beta_{-}\right)\left(\frac{m_{0}-M+i}{n}-\tau\right)+\alpha_{+}-\alpha_{-}\right)\right|$
$\geq\,$ $\frac{\left|\alpha_{+}-\alpha_{-}\right|}{d_{M}}\sum_{i=1}^{M}i-\frac{\left|\beta_{+}-\beta_{-}\right|}{d_{M}}\sum_{i=1}^{M}i\left(\frac{m_{0}-M+i}{n}-\tau\right)$
$\geq\,$ $\frac{3\left|\alpha_{+}-\alpha_{-}\right|}{2M+1}-\frac{\left|\beta_{+}-\beta_{-}\right|}{d_{M}n}\sum_{i=1}^{M}i(i+N)=\frac{3\left|\alpha_{+}-\alpha_{-}\right|}{2M+1}-\frac{\left|\beta_{+}-\beta_{-}\right|}{n}\left(1+\frac{3N}{2M+1}\right)>\frac{c}{n},$
for $n$ large enough. Set $Z_{m}\coloneqq(1/\sqrt{d_{M}})\sum_{i=1}^{M}i\varepsilon_{m-M+i}$ with $\varepsilon_{i}$ independently standard normal distributed. Next we define the event
$\mathcal{A}\coloneqq\left\{\max_{k<m\leq n}|Z_{m}|<\sqrt{8\log n},\max_{1\leq i\leq n}\left|f_{-}\left(\frac{i}{n}\right)-\hat{f}_{-}\left(\frac{i}{n}\right)\right|<\frac{c\left(2M+1\right)}{30n}\right\}.$ (19)
@@
$=\sqrt{\frac{3\log n}{b^{3}n^{2}\log n}}=\frac{c}{10n}.$ (20)
@@
$\left|\frac{1}{d_{M}}\sum_{i=1}^{M}i\left(f_{-}\left(\frac{m-M+i}{n}\right)-\hat{f}_{-}\left(\frac{m-M+i}{n}\right)\right)\right|<\frac{1}{d_{M}}\sum_{i=1}^{M}i\frac{c\left(2M+1\right)}{30n}\\ =\frac{6}{M\left(M+1\right)\left(2M+1\right)}\frac{c\left(2M+1\right)}{30n}\frac{M\left(M+1\right)}{2}=\frac{c}{10n}.$ (21)
@@
$\mathcal{A}\subset\left\{K_{m}<\frac{c}{5n}\text{ for }m<n\tau\text{ and }K_{m}\geq\frac{4c}{5n}\text{ for }m=\left\lceil n\tau\right\rceil+3N-1\right\}\subset\left\{n\tau\leq n\hat{\tau}_{K,n}\leq n\tau+3N\right\}$
@@
$\mathbb{P}_{\tau}\left(\max_{k<m\leq n}|Z_{m}|\geq\sqrt{8\log n}\right)\leq n\exp\left\{-4\log n\right\}=n^{-3}.$
@@
$\mathbb{P}_{\tau}\left(\max_{1\leq i\leq n}\left|f_{-}\left(\frac{i}{n}\right)-\hat{f}_{-}\left(\frac{i}{n}\right)\right|\geq\frac{c\left(2M+1\right)}{30n}\right)$
$\leq\,$ $\exp\left\{-\frac{c^{2}(2M+1)^{2}\left(c^{2}n^{2}-cn\right)}{300n^{2}(8cn+4)}\right\}+\exp\left\{-\frac{c^{2}(2M+1)^{2}(c^{3}n^{2}-c)}{28800n^{3}}\right\},$
$\asymp\,$ $\exp\left\{-n^{1/3}\log^{2/3}n\right\}\ll n^{-3}.$ (23)
@@
$\mathbb{P}_{\tau}\left(\mathcal{A}^{c}\right)\leq\mathbb{P}_{\tau}\left(\max_{1\leq m\leq M}|Z_{m}|\geq\sqrt{8\log M}\right)+\mathbb{P}_{\tau}\left(\max_{1\leq i\leq n}\left|f_{-}\left(\frac{i}{n}\right)-\hat{f}_{-}\left(\frac{i}{n}\right)\right|\geq\frac{c\left(2N+1\right)}{10n}\right)\leq n^{-3}.$
@@
$\max_{\theta_{0}\in\Theta^{K}}\mathbb{E}_{\tau}\left[\left(\frac{n^{1/3}}{\log^{1/3}n}\left(\tau_{K,n}-\tau\right)\right)^{2}\right]$
$=\,$ $\max_{\theta_{0}\in\Theta^{K}}\left(\mathbb{E}_{\tau}\left[\left(\frac{n^{1/3}}{\log^{1/3}n}\left(\tau_{K,n}-\tau\right)\right)^{2}\mathbb{I}(\mathcal{A})\right]+\mathbb{E}_{\tau}\left[\left(\frac{n^{1/3}}{\log^{1/3}n}\left(\tau_{K,n}-\tau\right)\right)^{2}\mathbb{I}(\mathcal{A}^{c})\right]\right)$
$\leq\,$ $\max_{\theta_{0}\in\Theta}\left(\mathbb{E}_{\tau}\left[\left(\frac{3N}{n^{2/3}\log^{1/3}n}\right)^{2}\mathbb{I}(\mathcal{A})\right]+\mathbb{E}_{\tau}\left[\left(\frac{n^{1/3}}{\log^{1/3}n}\right)^{2}\mathbb{I}(\mathcal{A}^{c})\right]\right)$
$\leq\,$ $\left(\frac{3N}{n^{2/3}\log^{1/3}n}\right)^{2}+\left(\frac{n^{1/3}}{\log^{1/3}n}\right)^{2}n^{-3}\leq 9b^{2}+1=9\cdot\left(\frac{300}{c^{2}}\right)^{2/3}+1,$